Run Coordination Meetings: Fall 2015 Run

From GlueXWiki
Jump to: navigation, search

Meeting General Agenda

Minutes of the September 22, 2015 meeting.

Minutes of the October 02, 2015 meeting.

Minutes of the October 24, 2015 meeting.

Minutes of the Dec. 1st, 2015 meeting.

Minutes of the Dec. 3rd, 2015 meeting.

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1]

2.) To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

4.) More information on connecting to bluejeans is available. The old ESNET and SeeVough Information is archived here.

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Dec. 8th 2015 meeting

  1. Report relevant updates from 7:50 and 8:00 daily MCC meetings:
    • Accelerator has steered beam through ARC 9 (4.5 passes) by end of owl shift last night. 1090MeV per linac (=9.9GeV)
    • Machine tuning is continuing. High trip rates (~15/h or more) required a RF Recovery. This is happening now. There was a problem for a power supply in ARC 9 (faulty choke). Accelerator has successfully switched it with choke from ARC 10. This was done in order to gain time and enable the delivery of beam with max. energy to Hall A. A new choke is being manufactured by the company (Danfysik, Denmark) and is expected by the end of this week or beginning of next week. Meanwhile, they are trying to repair the choke and it looks promising (it is working on the bench). Depending on this and on how long it takes to reach the important milestone of high power (900kw) 11 GeV in Hall A, we may start Saturday (tuning on Friday) with either a lower energy beam (for which the faulty choke can be used) or with the repaired choke. Or we would start at 12.05 GeV sometime around mid-next week, after the new choke arrived.
    • So, beam to Hall D is postponed to at least Saturday, Dec. 12th.
    • Link to accelerator short term schedule
    • Accelerator schedule through June 2017.
  2. Hot Check OUT Status 58.8% ready ; 3.4% checked ; 37.8% not ready
  3. Report on Work in the Hall:
    • Solenoid: Wrapping up repairs. Then, the cool down will begin. During this cool down, there will be times when control or restricted accesses will be needed (disruptions for MCC since any access for us bring down the gun).
    • Surveys
    • Amorphous Radiators
    • Radiation monitors Status of shielding?
    • Halo Counters : The one on the platform has been be put back. Others are good but there is a question regarding their alignment (orientation).
    • Microscope. Ready but need to do HCO.
    • Hodoscope checks.
    • Active Collimator - Placed on the 5mm hole.
    • Collimator On blocking position.
    • Profiler/intensity monitor - Profiler has been updated for faster acquisition. It working but a few checks remain.
    • Pair spectrometer. Ready.
    • Target Empty (i.e with H2 gas). A 1 cm plastic disk is in position at the end of the ST.
    • Start Counter There are 3 dead channels (most likely accidentally unplugged during ST motion). They will not be fixed since techs are busy with the solenoid and the ST will not be very useful for this run (target foil outside of the ST).
    • FDC
    • FCal
    • CDC One low voltage channel has its current readout dropping to zero and then coming back up. Some other channels display this sometimes too.
    • TOF
    • Transmission radiation detector
    • Slow controls. Mostly done. There are a couple of remaining things to do.
    • DAQ
    • Trigger integrating the tagger crates in the trigger.
    • Online monitoring
    • Tagger magnet, PS magnet, 2nd sweep magnet. NMR for tagger magnet has been fixed.
    • Goniometer
    • Polarimeter (converter is out of the way)
    • TAC
    • If you have work to do in the Hall, please notify Tom (Hall work coordinator) and if necessary, document it in the electronic logbook.
    • OSL need to be worn everywhere in Hall D.
  4. Fall 2015 Commissioning Plan